Be original amd used your own word
1. Go to Yahoo! Finance or another source and pick a stock to track for the last year. Discuss the performance of the stock over the last year and offer your thoughts being specific about why the stock has done well or not.


Add your refrence
Answer #1

Source: Yahoo Finance

Stock: Alphabet (Google)

Stock Price and Return:

  1. 12th Nov 2018 - 1,061.49
  2. 13th Nov 2019 - 1,299.36
  3. Return - 1,299.36/1,061.49 = 22.41%

Market Return:

  1. 12th Nov 2018 - 25,413.22
  2. 13th Nov 2019 - 27,686.50
  3. Return - 27,686.50/25,413.22 = 8.95%

Google has outperformed the index in last 12 months as Dow gave returns of 8.95% in last one year whereas Google gave returns of 22.41%.

Google Revenue (in millions) on Dec 2018 Quarter End - 39,276

Google Revenue (in millions) on Sept 2019 Quarter End - 40,499

Operating Income (in millions) on Dec 2018 Quarter End - 8,203

Operating Income (in millions) on Sept 2019 Quarter End - 9,117

If we further analyze results few facts.

Globally economy is under pressure and living with fear of recession, however over a period of time Google has been evolved as recession free business model. Gradually contribution from Ad Revenue is increasing irrespective of Economy performance. Digital spending for marketing activities is increasing like never before and Google is benefiting the most out of this change.

If we compare the revenue performance from Dec 2018 to Sept 2019, there is 3.11% increase in revenue but if we look at operating income there is jump of 14.39%. This shows operational efficiency increasing drastically.

Operational efficiency can also be calculated by Operating Income / Revenue

Operational Efficiency Dec 2018 = 8,023/39,279 = 20.43%

Operational Efficiency Sept 2019 = 9,177/40,499 = 22.66%

This shows company is becoming better at operational front any any increase in revenue would impact better at profit front.

Thanks